An auditor reviewed 25 oral surgery insurance claims from a particular surgical office, determining that the mean out-of-pocket patient billing above the reimbursed amount was $275.66 with a standard deviation of $78.11. (a) At the 5 percent level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $250 out-of-pocket? State your hypotheses and decision rule. (b) Is this a close decision?
Answer by stanbon(29602)

An auditor reviewed 25 oral surgery insurance claims from a particular surgical office, determining that the mean out-of-pocket patient billing above the reimbursed amount was $275.66 with a standard deviation of $78.11.
---------------------

(a) At the 5 percent level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $250 out-of-pocket? State your hypotheses and decision rule.
---
Ho: u <= 250
Ha: u > 250
----------------
Critical value for a right-tailed test with alpha = 5%, df=24: t=1.711
--------------------
Test Statistic:
t(275.66) = (275.66-250)/[78.11/sqrt(25)] = 1.643
p-value = P(t > 1.643) = 0.05676
----------------
Conclusion: Since the p-value is greater than 5%, Fail to reject Ho.
-------------------
(b) Is this a close decision?
Yes, very close as the p-value is "close" to 5%.
=======================================================
Cheers,
Stan H.